# Supersingular curves via the Shimura–Taniyama method

### v1

@@ 2.5. Decomposition from the Kani–Rosen theorem
(2.3) $\mathrm{Jac}(X)^{t-1}\times\mathrm{Jac}(X/G)^{n_{G}}\sim\prod_{i=1}^{t}\mathrm{Jac}(X/H_{i})^{n_{i}}.$
Suppose $G\simeq(\mathbf{Z}/2\mathbf{Z})^{2}$ and let $H_{1},H_{2},H_{3}\subset G$ be the three subgroups of order two. Then (2.3) simplifies to $\mathrm{Jac}(X)\times\mathrm{Jac}(X/G)^{2}\sim\prod_{i=1}^{3}\mathrm{Jac}(X/H_{i})$. Suppose $G\simeq(\mathbf{Z}/3\mathbf{Z})^{2}$ and let $H_{1},H_{2},H_{3},H_{4}\subset G$ be the four subgroups of order three. Then (2.3) simplifies to $\mathrm{Jac}(X)\times\mathrm{Jac}(X/G)^{3}\sim\prod_{i=1}^{4}\mathrm{Jac}(X/H_{i})$.
@@ 2.6. Enumerating and analyzing abelian covers
Using SageMath, we automated the search for ramification data producing abelian covers $\phi:X\to\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$ whose genus lies in a specified range. We outline the approach, which depends on whether $s=1$ and $s>1$.
When $s=1$, then $\phi$ is cyclic and the program determines possible inertia types. Without loss of generality, we identified the generator of $I_{0}$ with $1\in G$ and so $a_{0}=1$. Then the program analyzes the congruence classes of primes $p$ for which the reduction of $X$ modulo $p$ is supersingular, using the Shimura–Taniyama formula as in Section 2.4.
For each case when $s>1$, Lemma 2.6 implies that $G$ is non-cyclic and that $G\simeq(\mathbf{Z}/c_{0}\mathbf{Z})\times(\mathbf{Z}/s\mathbf{Z})$. For every $z$, we determined the possible inertia types $a$. Without loss of generality, we identified the generator of $I_{0}$ with $a_{0}=(1,0)$. In the small number of non-cyclic cases with $5\leq g\leq 10$, we determined by hand a decomposition of $\mathbf{Jac}(X)$ into Jacobians of curves that are cyclic covers of $\mathbf{P}^{1}$ branched at $B$, using the Kani–Rosen method as in Section 2.5. This allowed us to find the supersingular primes.
In Section 3, we give a complete analysis of the genus $5$ situation to illustrate the methods. In particular, abelian covers of $\mathbf{P}^{1}$ branched at three points do not produce supersingular curves of genus $5$ for every prime $p\equiv 3\bmod 4$. Section 4 contains the results of our program [BP25] for cyclic cases when $6\leq g\leq 10$. Section 5 treats the non-cyclic cases when $6\leq g\leq 10$.
@@ 3. Refining Ekedahl’s Results for genus five
Suppose $\phi:X\to\mathbf{P}^{1}$ is a tamely ramified abelian cover branched at $B=\{0,1,\infty\}$. In [Eke87, page 173], Ekedahl enumerates ways $\phi$ can ramify so that $X$ has genus $5$. There are seven non-equivalent cases, as confirmed by our program, which are listed in Propositions 3.1, 3.3, and 3.6.
Ekedahl studies the situation when $X$ is superspecial, meaning that $\mathrm{Jac}(X)$ is isomorphic to a product of supersingular elliptic curves. The superspecial condition is equivalent to the Frobenius operator being trivial on $H^{1}(X,\mathcal{O}_{X})$. For each of the seven cases, Ekedahl determines the primes $p$ for which $X$ is superspecial, thus proving that there exists a superspecial smooth curve of genus $5$ which is an abelian cover of $\mathbf{P}^{1}$ branched at $3$ points if and only if $p\equiv-1\bmod 11,12,15,20,8$ or $p\equiv 11\bmod 15$.
@@ 3.1. Cyclic cases for genus five
If $\phi:X\to\mathbf{P}^{1}$ is a cyclic cover branched at $3$ points and $X$ has genus $5$, then the ramification type $z$ for $\phi$ is one of the following:
  1. (1)
    $z=[11,11,11,1]$. Then $X$ is supersingular if and only if $p$ is a QNR modulo $11$.
  2. (2)
    $z=[15,15,3,1]$. Then $X$ is supersingular if and only if $p\equiv-1,11\bmod 15$.
  3. (3)
    $z=[22,11,2,1]$. Then $X$ is supersingular if and only if $p$ is a QNR modulo $11$.
  4. (4)
    $z=[12,12,6,1]$. Then $X$ is supersingular if and only if $p\equiv-1\bmod 12$.
  5. (5)
    $z=[20,20,2,1]$. Then $X$ is supersingular if and only if $p\equiv-1,11\bmod 20$.
@@
Recall that $R$ denotes the set of orbits of ×$\times p$ on $\mathbf{Z}/m\mathbf{Z}$. In each situation, if $p^{i}\equiv-1\bmod m$ for some $i$, then each orbit is self-dual and so $X$ is supersingular. When $m$ is composite, there can be additional cases where $\alpha_{r}/\#r=1/2$ for each $r\in R$. We give some additional details in each case.
@@
    Note that $a\approx(1,k,-(1+k))$ for some $1\leq k\leq 9$. The result is proven as one case of [LMPT19b, Theorem 5.1].
@@
    Then $a\approx(1,9,10)$ and $S_{1}=\{1,3,5,7,9\}$. When $p\equiv 11\bmod 20$, then #$\#r=2$ for each $r\in R$ and each orbit contains exactly one element of $S_{1}$. By the Shimura–Taniyama formula, $X$ is supersingular (but not superspecial). ∎
@@ 3.2. Non-cyclic cases for genus five
Suppose $\phi:X\to\mathbf{P}^{1}$ is an abelian non-cyclic cover branched at $B=\{0,1,\infty\}$ and $X$ has genus $5$. Ekedahl shows there are two options for the ramification type: either (6) $z=[8,8,4,2]$ or (7) $z=[12,12,2,2]$. In both cases, we determine the congruence conditions on $p$ for which $X$ is supersingular.
@@ 3.2.1. Case (6)
When $z=[8,8,4,2]$, then:
  1. (1)
    the quotient of $X$ by $H$ is isomorphic to $E_{1728}$;
@@
    Note that $G/H\simeq\mathbf{Z}/4\mathbf{Z}$. So $\psi:X/H\to{\mathbf{P}}^{1}$ is a $\mathbf{Z}/4\mathbf{Z}$-cover branched at $B$. The intersection of $H$ with each of $I_{0}$, $I_{1}$, and $I_{\infty}$ has order $2$. By the Riemann–Hurwitz formula, $X/H$ has genus $1$. Since $\psi$ is totally ramified above 00 and $\infty$, this implies that $X/H$ is isomorphic to $E_{1728}$.
@@
    Let $H_{1}:=\langle(4,0)\rangle$. Since $(4,0)$ is in the inertia group at every point of $X$ above 00, $1$, and $\infty$, the double cover $X\to X/{H_{1}}$ is branched at $8$ points. By the Riemann–Hurwitz formula, $X/{H_{1}}$ is an elliptic curve, which we denote $E^{\prime}$. The double cover $E^{\prime}\to E_{1728}$ is unramified and is thus a $2$-isogeny.
    Let $W$ be the quotient of $X$ by $H_{2}:=\langle(0,1)\rangle$. Since $H_{2}\cap I_{b}$ is trivial for each $b\in B$, the double cover $X\to W$ is unramified. By the Riemann–Hurwitz formula, $W$ has genus $3$. The same holds for the quotient $W^{\prime}$ of $X$ by $H_{3}:=\langle(4,1)\rangle$.
  3. (3)
    By Theorem 2.12, ${\mathbf{Jac}}(X)\times{\mathbf{Jac}}(E_{1728})^{2}\sim{\mathbf{Jac}}(E^{\prime})\times{\mathbf{Jac}}(W)\times{\mathbf{Jac}}(W^{\prime})$. The result follows since the category of abelian varieties up to isogeny is semi-simple, and ${\mathbf{Jac}}(E_{1728})\sim{\mathbf{Jac}}(E^{\prime})$. ∎
@@
Suppose $z=[8,8,4,2]$. Then $X$ is supersingular if and only if $p\equiv 7\bmod 8$. If $p\equiv 3\bmod 8$, then ${\mathbf{Jac}}(X)$ has Newton polygon $(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}^{3}$.
@@
By Lemma 3.2(i), $E_{1728}$ is a quotient of $X$. We restrict to the case $p\equiv 3\bmod 4$ which is equivalent to $E_{1728}\sim E^{\prime}$ being supersingular. In this case, $X$ is supersingular if and only if $W$ and $W^{\prime}$ are both supersingular, by Lemma 3.2(iii).
@@
If $p\equiv 7\bmod 8$, then every orbit of ×$\times p$ on $\{1\leq j<8\}$ is self-dual, and so both $W$ and $W^{\prime}$ are supersingular by the Shimura–Taniyama formula. Suppose $p\equiv 3\bmod 8$. By Example 2.10, the curve $W^{\prime}$ is supersingular since $\overline{a}^{\prime}\approx(1,5,2)$, while the curve $W$ has Newton polygon $(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}$ since $\overline{a}=(1,1,6)$. Thus ${\mathbf{Jac}}(X)\times{\mathbf{Jac}}(E_{1728})$ has two slopes of 00 and $1$, and $X$ has Newton polygon $(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}^{3}$. ∎
@@ 3.2.2. Case (7)
    The curve $Z$ has an action by $G/\langle(6,1)\rangle\simeq\mathbf{Z}/12\mathbf{Z}$, which is generated by $w=(1,1)$. The inertia type $a=((1,0),(11,1),(0,1))$ reduces to $\overline{a}=(7w,11w,6w)\approx(5,1,6)$. There is an order $3$ automorphism $\sigma$ on $Z$ that fixes two points. The quotient of $Z$ by $\sigma$ is an elliptic curve with an automorphism of order $4$ that fixes two points. Thus $Z$ is a $\mathbf{Z}/3\mathbf{Z}$-cover of $E_{1728}$.
@@
    The curve $W$ admits an action of $\mathbf{Z}/6\mathbf{Z}\times\mathbf{Z}/2\mathbf{Z}$. Each of the automorphisms $(3,0)$ and $(3,1)$ fixes exactly two points of $W$, so $W$ admits two double covers of an elliptic curve. In both cases, this elliptic curve is $E_{0}$ because it has an automorphism of order $3$ which fixes three points. Since $(3,0)$ and $(3,1)$ generate a Klein-four group, ${\mathbf{Jac}}(W)\sim E_{0}^{2}$ by Theorem 2.12. ∎
@@
Suppose $z=[12,12,2,2]$.
Then $X$ is supersingular if and only if $p\equiv 11\bmod 12$.
If $p\equiv 5\bmod 12$, then ${\mathbf{Jac}}(X)$ has Newton polygon $(G_{0,1}\oplus G_{1,0})^{3}\oplus G_{1,1}^{2}$.
If $p\equiv 7\bmod 12$, then ${\mathbf{Jac}}(X)$ has Newton polygon $(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}^{3}$.
@@
Suppose $p\equiv 11\bmod 12$. Then $E_{0}$ is supersingular. By Lemma 3.5, ${\mathbf{Jac}}(X)\sim{\mathbf{Jac}}(Z)\times E_{0}^{2}$. The curve $Z$ is supersingular by the Shimura–Taniyama formula since the orbits of ×$\times p$ on $\mathbf{Z}/12\mathbf{Z}$ are self-dual. This completes the proof of the first claim.
The $\mathbf{Z}/12\mathbf{Z}$-cover $Z\to\mathbf{P}^{1}$ has inertia type $(5,1,6)$. We compute that the eigenspaces of $H^{0}(Z,\Omega^{1}_{Z})$ of positive dimension are indexed by $S_{1}=\{1,3,5\}$. When $p\equiv 5\bmod 12$, then $E_{0}$ is supersingular. Also the orbits of ×$\times p$ are $\{1,5\}$, $\{7,11\}$, $\{3\}$ and $\{9\}$. By the Shimura–Taniyama formula, the Newton polygon of $Z$ is $(G_{0,1}\oplus G_{1,0})^{3}$. When $p\equiv 7\bmod 12$, then $E_{0}$ is ordinary. Also the orbits of ×$\times p$ are $\{1,7\}$, $\{5,11\}$, and $\{3,9\}$. By the Shimura–Taniyama formula, $Z$ is supersingular. ∎

### v2

@@ 2.6. Decomposition from the Kani–Rosen theorem
(5) $\mathrm{Jac}(X)^{t-1}\times\mathrm{Jac}(X/G)^{n_{G}}\sim\prod_{i=1}^{t}\mathrm{Jac}(X/H_{i})^{n_{i}}.$
Example 2.13.
Suppose $G\simeq(\mathbf{Z}/2\mathbf{Z})^{2}$ and let $H_{1},H_{2},H_{3}\subset G$ be the three subgroups of order two. Then (5) simplifies to $\mathrm{Jac}(X)\times\mathrm{Jac}(X/G)^{2}\sim\prod_{i=1}^{3}\mathrm{Jac}(X/H_{i})$. Suppose $G\simeq(\mathbf{Z}/3\mathbf{Z})^{2}$ and let $H_{1},H_{2},H_{3},H_{4}\subset G$ be the four subgroups of order three. Then (5) simplifies to $\mathrm{Jac}(X)\times\mathrm{Jac}(X/G)^{3}\sim\prod_{i=1}^{4}\mathrm{Jac}(X/H_{i})$.
@@ 2.7. Enumerating and analyzing abelian covers
Using SageMath, we automated the search for abelian covers $\phi:X\to\mathbf{P}^{1}$ over $\mathbf{Q}$ branched at $B=\{0,1,\infty\}$ whose genus lies in a specified range. We then partially automated the determination of the set of primes $p$ for which the reduction of $X$ modulo $p$ is supersingular. We outline the approach here: the code is on github [BP25].
We begin by listing all possible $e_{0,1},e_{0,\infty},e_{1,\infty},r,s$ which satisfy equation (2) for $g$ in the specified range. These determine the potential ramification types of covers.
For a potential ramification type with $s=1$, the cover $\phi$ is cyclic and $G=\mathbf{Z}/m\mathbf{Z}$. We look for inertia types compatible with the ramification type. Without loss of generality, $a_{0}=1$, and we search for $a_{1}$ and $a_{\infty}$ with $|a_{1}|=c_{1}$, $|a_{\infty}|=c_{\infty}$, and $a_{0}+a_{1}+a_{\infty}=0\mod{m}$. Then (3) gives a cover with that inertia type. Finally, the program determines the congruence classes of primes $p$ for which the reduction of $X$ modulo $p$ is supersingular, using the Shimura–Taniyama formula found in Theorem 2.8.
When $s>1$, Lemma 2.11 implies that $\phi$ is non-cyclic and that $G\simeq(\mathbf{Z}/c_{0}\mathbf{Z})\times(\mathbf{Z}/s\mathbf{Z})$. Again, we look for inertia types compatible with the ramification type. Without loss of generality, we identify the potential generator of $I_{0}$ with $a_{0}=(1,0)$ and search for $a_{1}$ and $a_{\infty}$ satisfying the constraints listed after Definition 2.4. Note that an inertia type realizing a potential ramification type may not exist, in which case there is no cover $\phi$ realizing that ramification type; see Remark 5.1 for examples.
In the small number of non-cyclic cases with $5\leq g\leq 10$, for a curve $X$ with inertia type $a$, we determine by hand a decomposition of $\mathbf{Jac}(X)$ into Jacobians of curves that are cyclic covers of $\mathbf{P}^{1}$ branched at $B$, using the Kani–Rosen method as in Section 2.6. The program then gives us the primes for which the reduction of $X$ is supersingular.
In Section 3, we give a complete analysis of the genus $5$ situation to illustrate the methods. Section 4 contains the results of our program for cyclic cases when $6\leq g\leq 10$. Section 5 treats the non-cyclic cases when $6\leq g\leq 10$.
@@ 3. Refining Ekedahl’s Results for genus five
Suppose $\phi:X\to\mathbf{P}^{1}$ is a tamely ramified abelian cover over $k$ branched at $B=\{0,1,\infty\}$. In [Eke87, page 173], Ekedahl enumerates ways $\phi$ can ramify so that $X$ has genus $5$. There are seven non-equivalent cases, as confirmed by our program, which are listed in Propositions 3.1, 3.3, and 3.6.
Ekedahl studies the situation when $X$ is superspecial, meaning that $\mathrm{Jac}(X)$ is isomorphic to a product of supersingular elliptic curves. The superspecial condition is equivalent to the Frobenius operator being trivial on $H^{1}(X,\mathcal{O}_{X})$. For each of the seven cases, Ekedahl determines the primes $p$ for which $X$ is superspecial, thus proving that there exists a superspecial smooth curve of genus $5$ which is an abelian cover of $\mathbf{P}^{1}$ branched at $3$ points if and only if $p\equiv-1\bmod 8,11,12,15,20$ or $p\equiv 11\bmod 15$.
@@ 3.1. Cyclic cases for genus five
If $\phi:X\to\mathbf{P}^{1}$ is a cyclic cover branched at $3$ points and $X$ has genus $5$, then its ramification type $z$ and set of supersingular primes are one of:
Case z supersingular primes
$1$ $[11,11,11,1]$ $p\equiv\text{QNR}\bmod 11$
$2$ $[15,15,3,1]$ $p\equiv-1,11\bmod 15$
$3$ $[22,11,2,1]$ $p\equiv\text{QNR}\bmod 11$
$4$ $[12,12,6,1]$ $p\equiv-1\bmod 12$
$5$ $[20,20,2,1]$ $p\equiv-1,11\bmod 20$
@@
Recall that $R$ denotes the set of orbits of ×$\times p$ on $\mathbf{Z}/m\mathbf{Z}$. If $p^{i}\equiv-1\bmod m$ for some $i$, then each orbit is self-dual and so $X$ is supersingular. When $m$ is composite, there can be additional cases where $\alpha_{r}/\#r=1/2$ for each $r\in R$. Here are some details.
@@
    This is one case of [LMPT19b, Theorem 5.1].
@@
    Then $a\approx(1,9,10)$ and $S_{1}=\{1,3,5,7,9\}$. When $p\equiv 11\bmod 20$, then #$\#r=2$ for each $r\in R$ and each orbit contains exactly one element of $S_{1}$. By the Shimura–Taniyama formula, $X$ is supersingular (but not superspecial).
@@ 3.2. Non-cyclic cases for genus five
Suppose $\phi:X\to\mathbf{P}^{1}$ is an abelian non-cyclic cover over $k$ branched at $B=\{0,1,\infty\}$. When $X$ has genus $5$, Ekedahl shows there are two options for the ramification type: either (6) $z=[8,8,4,2]$ or (7) $z=[12,12,2,2]$.
@@ 3.2.1. Case (6)
When $z=[8,8,4,2]$, then $X/H$ is isomorphic to $E_{1728}$. Also:
@@
Note that $G/H\simeq\mathbf{Z}/4\mathbf{Z}$. So $\psi:X/H\to{\mathbf{P}}^{1}$ is a $\mathbf{Z}/4\mathbf{Z}$-cover branched at $B$. The intersection of $H$ with each of $I_{0}$, $I_{1}$, and $I_{\infty}$ has order $2$, so $X/H$ has genus $1$. Since $\psi$ is totally ramified above 00 and $1$, this implies that $X/H\cong E_{1728}$.
@@
    Let $H_{1}:=\langle(4,0)\rangle$. Since $(4,0)$ is in the inertia group at every point of $X$ above 00, $1$, and $\infty$, the double cover $X\to X/{H_{1}}$ is branched at $8$ points. So $X/{H_{1}}$ is an elliptic curve, which we denote $E^{\prime}$. The double cover $E^{\prime}\to E_{1728}$ is unramified and is thus a $2$-isogeny. Let $W$ be the quotient of $X$ by $H_{2}:=\langle(0,1)\rangle$. Since $H_{2}\cap I_{b}$ is trivial for each $b\in B$, the double cover $X\to W$ is unramified and $W$ has genus $3$. The same holds for the quotient $W^{\prime}$ of $X$ by $H_{3}:=\langle(4,1)\rangle$.
@@
    By Theorem 2.12, ${\mathbf{Jac}}(X)\times{\mathbf{Jac}}(E_{1728})^{2}\sim{\mathbf{Jac}}(E^{\prime})\times{\mathbf{Jac}}(W)\times{\mathbf{Jac}}(W^{\prime})$. The result follows since the category of abelian varieties up to isogeny is semi-simple, and ${\mathbf{Jac}}(E_{1728})\sim{\mathbf{Jac}}(E^{\prime})$.
@@
Suppose $z=[8,8,4,2]$. Then $X$ is supersingular if and only if $p\equiv 7\bmod 8$. If $p\equiv 3\bmod 8$, then $\nu({\mathbf{Jac}}(X))=(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}^{3}$.
@@
By Lemma 3.2, $E_{1728}$ is a quotient of $X$. We restrict to the case $p\equiv 3\bmod 4$ which is equivalent to $E_{1728}\sim E^{\prime}$ being supersingular. In this case, $X$ is supersingular if and only if $W$ and $W^{\prime}$ are both supersingular, by Lemma 3.2.
@@
If $p\equiv 7\bmod 8$, then every orbit of ×$\times p$ on $\{1\leq j<8\}$ is self-dual, and so both $W$ and $W^{\prime}$ are supersingular by Theorem 2.8. Suppose $p\equiv 3\bmod 8$. By Example 2.9, the curve $W^{\prime}$ is supersingular since $\overline{a}^{\prime}\approx(1,5,2)$, while $\nu(\mathbf{Jac}(W))=(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}$ since $\overline{a}=(1,1,6)$. Thus ${\mathbf{Jac}}(X)\times{\mathbf{Jac}}(E_{1728})$ has two slopes of 00 and $1$, and $\nu({\mathbf{Jac}}(X))=(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}^{3}$. ∎
@@ 3.2.2. Case (7)
    The curve $Z$ has an action by $G/\langle(6,1)\rangle\simeq\mathbf{Z}/12\mathbf{Z}$, which is generated by $w=(1,1)$. The inertia type $a$ reduces to $\overline{a}=(7w,11w,6w)\approx(5,1,6)$. The curve $Z$ has an order $3$ automorphism $\sigma$ that fixes two points and $Z/\langle\sigma\rangle$ is an elliptic curve with an automorphism of order $4$ that fixes two points. Thus $Z$ is a $\mathbf{Z}/3\mathbf{Z}$-cover of $E_{1728}$.
@@
    The curve $W$ admits an action of $\mathbf{Z}/6\mathbf{Z}\times\mathbf{Z}/2\mathbf{Z}$. The automorphisms $(3,0)$ and $(3,1)$ each fix two points of $W$, so $W$ admits two double covers of an elliptic curve, which must be $E_{0}$ because it has an automorphism of order $3$ which fixes three points. Since $(3,0)$ and $(3,1)$ generate a Klein-four group, ${\mathbf{Jac}}(W)\sim E_{0}^{2}$ by Theorem 2.12.
@@
Suppose $z=[12,12,2,2]$. Then $X$ is supersingular if and only if $p\equiv 11\bmod 12$. If $p\equiv 5\bmod 12$, then $\nu({\mathbf{Jac}}(X))=(G_{0,1}\oplus G_{1,0})^{3}\oplus G_{1,1}^{2}$. If $p\equiv 7\bmod 12$, then $\nu({\mathbf{Jac}}(X))=(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}^{3}$.
@@
Suppose $p\equiv 11\bmod 12$. Then $E_{0}$ is supersingular. By Lemma 3.5, ${\mathbf{Jac}}(X)\sim{\mathbf{Jac}}(Z)\times E_{0}^{2}$. The curve $Z$ is supersingular by Theorem 2.8, since the orbits of ×$\times p$ on $\mathbf{Z}/12\mathbf{Z}$ are self-dual. This completes the proof of the first claim.
The $\mathbf{Z}/12\mathbf{Z}$-cover $Z\to\mathbf{P}^{1}$ has inertia type $(5,1,6)$. We compute that the eigenspaces of $H^{0}(Z,\Omega^{1}_{Z})$ of positive dimension are indexed by $S_{1}=\{1,3,5\}$. When $p\equiv 5\bmod 12$, then $E_{0}$ is supersingular. Also, the orbits of ×$\times p$ are $\{1,5\}$, $\{7,11\}$, $\{3\}$ and $\{9\}$. By Theorem 2.8, $\nu(\mathbf{Jac}(Z))=(G_{0,1}\oplus G_{1,0})^{3}$. When $p\equiv 7\bmod 12$, then $E_{0}$ is ordinary. Also, the orbits of ×$\times p$ are $\{1,7\}$, $\{5,11\}$, and $\{3,9\}$. By Theorem 2.8, $Z$ is supersingular. ∎